# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 3.4 Anomalous supercharged nuclei consisting dark matter
In 1981 Cahn and Glashow Cahn and Glashow (1981) assumed existence of superheavy charged dark matter particles, cf. also De Rujula et al. (1990). In 2005 Glashow Glashow (2005) conjectured that more specifically the dark matter may consist of the U tera-quarks and O-UUU (charge 2-2) baryons, and in 2006 Khlopov Khlopov (2006) studied OOHe-hidden dark matter with $m_{O}\sim\!1$ TeV. Reference Gani et al. (2019) suggested existence of stable $O$-nuclearites of arbitrary size (up to $A\sim 10^{57}$), if there exist charged dark matter stable superheavy particles. In this case the proton charge of the nucleus is compensated by the heavy $O$-particles. Due to the large mass of the $O$ particles they yield a negligible contribution to the kinetic energy. Therefore, the ordinary isospin-symmetric matter at density $n\simeq n_{0}$ with embedded $O$-particles compensating the charge would be absolutely stable. Similarly to estimates (17), (20) performed above for supercharged nuclei, in case of the nuclear–$O$ matter one may present the energy as
${\cal{E}}\simeq\mathcal{E}_{0}\times A-\int d^{3}x(n_{p}-2n_{O})V\,.$ (24)
@@
At present there exist various anomalous events, some of which may have a relation to the discussed above scalar fields, $\pi^{-}$ condensate, quark or dark matter nuggets, see Zhitnitsky and Maroudas (2025) and references therein and a discussion in Sec. 8 below.
4 Superheavy nuclear objects
4.1 $\pi^{\pm}$ condensate nuclear systems
@@
The phenomenon of a pion condensation was proposed and investigated in Refs. Migdal (1971, 1973); Scalapino (1972); Sawyer (1972), see also Migdal (1978); Migdal et al. (1990); Voskresensky (1993) and references therein. The pion-nucleon interaction is strongly attractive in the p-wave. Therefore, from the first works one expected that the pion condensate will be formed with a finite momentum $k\neq 0$ when the nucleon density exceeds some critical density $n_{c}^{p}$ Migdal (1973a, b); Baym (1973); Migdal et al. (1974); Campbell et al. (1975a, b); Baym et al. (1975); Dautry and Nyman (1979); Takatsuka et al. (1978). The $\pi N$ p-wave attraction is so strong that initially one anticipated that $n_{c}^{p}$ might be smaller than $n_{0}$. Soon, after performed analysis of the atomic nucleus data, it became clear that $n_{c}^{p}>n_{0}$ because of a stronger repulsive effect of the short-range nucleon-nucleon correlations, see  Migdal (1978); Migdal et al. (1990) and references therein.
@@
Main contributions to the pion polarization operator come from the nucleon-nucleon hole and $\Delta$-nucleon hole terms both corrected by the short-range correlations, for more details see Migdal et al. (1990); Voskresensky (2023). The retarded pion propagator is
@@
$-[D_{\pi}^{R}(\mu_{\pi},\vec{k})]^{-1}=\widetilde{\omega}^{2}(\mu_{\pi},k^{2})=m^{2}_{\pi}+k^{2}+\Re\Pi^{R}(\mu_{\pi},k^{2})-\mu^{2}_{\pi}\,$, (26)
$\Re\Pi(\omega,k)$ is the real part of the pion polarization operator, the imaginary part $\Im\Pi(\mu_{\pi},k)=0$, $f_{\pi NN}\approx m_{\pi}^{-1}$ is the $\pi NN$ coupling constant, and $\Gamma(n)$ is the factor taking into account the $NN$ correlations, $\Gamma(n\to 0)\to 1$, $\mu_{\pi}$ is the pion chemical potential. In case of the isospin symmetric matter $\mu_{\pi^{\pm}}=0$ in accordance with reactions $p\leftrightarrow n+\pi^{+}$, $n\leftrightarrow p+\pi^{-}$; $\mu_{\pi^{0}}=0$ for the case of arbitrary isospin composition in accordance with reactions $N\leftrightarrow N+\pi^{0}$, $N=(n,p)$.
Spectra of pions in the isospin-symmetric and purely neutron matter were demonstrated in Migdal et al. (1974). In the latter case for $\pi^{\pm}$ there are the $\pi^{+}$ and $\pi^{-}$ branches, $\omega^{\pi^{+}}(k)$, $\omega^{\pi^{-}}(k)$ (corresponding to $\omega\to m_{\pi}$ for $n\to 0$) and the higher lying $\Delta$ branches. It proves to be that up to large densities $\omega^{\pi^{-}}(k)>\mu_{e}$ Migdal (1973a); Migdal et al. (1974) and the reaction $N+e\to N+\pi^{-}+\nu$ is forbidden. If pion were free, $\omega^{2}=m^{2}_{\pi}+k^{2}$, such a reaction were allowed for $\mu_{e}(n)>m_{\pi}$ resulting in efficient cooling of the neutron stars heaving such a densities in their interiors. For $n>n_{c}^{+}$ ($n_{c}^{+}<n_{0}$ according to evaluations of Migdal et al. (1974)) there appears extra branch with $\omega^{\pi^{+}}_{s}(k)<0$ corresponding to appearance of the so called $\pi^{+}_{s}$ condensate. Accordingly, for $n>n_{c}^{+}$ the reaction $p\to n+\pi^{+}$ becomes possible resulting in the production of $\pi^{+}$’s occupying this branch. For $n=n_{c}^{\pm}>n_{c}^{+}$ the branches $\omega^{\pi^{+}}_{s}(k)$ and $\omega^{\pi^{-}}(k)$ coincide at $k=k_{0}\neq 0$, where $\omega^{\pi^{+}}_{s}(k_{0})+\omega^{-}_{\pi}(k_{0})=0$. Thereby for $n>n_{c}^{\pm}$ the reactions $N\leftrightarrow N+\pi^{+}+\pi^{-}$ become energetically favorable and there appears $\pi^{\pm}$ condensate. Appearance of $\pi^{\pm}$ condensation for $n>n_{c}^{\pm}$ results in the efficient neutrino cooling processes, e.g. via $N+\pi_{cond}^{-}\to N+e+\bar{\nu}$ reaction.
The quantity
$\widetilde{\omega}^{2}_{0}=\min_{k}\{\widetilde{\omega}^{2}(\mu_{\pi},k)\}=\widetilde{\omega}^{2}(\mu_{\pi},k_{0})\,$, (27)
where the minimum is realized at a density dependent momentum $k_{0}(n)$, has the meaning of the squared effective pion gap, see Migdal et al. (1990) and references therein. We notice that the in-medium $\pi^{-}$ plays a role somewhat similar to the light negatively charged boson considered in the previous section, the boson squared mass, $m_{b}^{2}$, should be replaced by the squared effective pion gap.
At densities $n\sim n_{c}^{p}$ the pion gap behaves like $\widetilde{\omega}^{2}_{0}\sim\beta(n_{c}^{p}-n)$, with a constant $\beta(\mu_{\pi},k_{0})>0$ (at the neglect of the quantum fluctuations) and the change in the sign of the gap indicates the onset of instability with respect to the pion condensation with the momentum $k_{0}$. The critical point $n_{c}^{p}$ is the same for $\pi^{\pm},\pi^{0}$ in isospin symmetric matter. In the neutron star matter $n_{c}^{p}$ coincides with $n_{c}^{\pm}$ for charged pions and it is in general different for $\pi^{0}$.
In the isospin-symmetric matter the momentum dependence of the pion gap, $\widetilde{\omega}_{0}^{2}(k^{2})=\widetilde{\omega}^{2}(0,k^{2})$ develops a minimum at $k=k_{0}\neq 0$ at densities $n>n_{c}^{(1)}\simeq(0.5-0.7)\,n_{0}$, see Voskresensky (1993), and in its vicinity
@@
where $\gamma\sim 1$, and $\widetilde{\omega}^{2}_{0}$ changes sign at $n=n_{c}^{p}$. Considering the pion condensation in a finite system for $n>n_{c}^{p}>n_{0}$, the spatial structure of the classical pion field is governed by the equation
@@
which replaces Eq. (18). Nuclear systems obeying the roton-like spectra have been extensively studied, see Voskresensky and Anisimov (1980); Migdal et al. (1990); Voskresensky (2023) and references therein. From a recent time some authors became to name such a spectrum the “moat” spectrum.
@@
At large densities the baryon subsystem is essentially rearranged. In the approach exploiting the chiral symmetry  Campbell et al. (1975a, b); Baym et al. (1975) it was shown that instead of two different Fermi seas for protons and neutrons, a common Fermi sea for baryon quasiparticles, being superpositions of proton neutron and $\Delta$ isobar states mixed due to interactions with the developed pion condensate, is formed. At densities much larger than $n_{c}^{p}$ the squared amplitude of the pion condensate reaches its maximum value $f_{\pi}^{2}/2$ for $\pi^{\pm}$, where $f_{\pi}\approx 92$ MeV is the pion weak decay constant and the electric charge density of the baryon subsystem becomes equal to $n/2$. Taking into account these effects, and electron and muon populations, the energy density acquires the form Voskresensky et al. (1977)
${E}_{B+\pi}$ $={E}_{B}+{E}_{\pi}+{E}_{e}+{E}_{\mu}\,$,
@@
${E}_{\pi}$ $=\frac{V^{2}f^{2}_{\pi}}{2}+\frac{1}{3}(m^{*}_{\Delta}-m^{*}_{N})\,n-\frac{81}{50}f^{2}_{\pi NN}(1-g^{\prime})n^{2}\,$, (28)
where $E_{B}$ is the energy density of the baryonic subsystem, $g^{\prime}$ is the isospin Landau-Migdal parameter, and $m_{\Delta}^{*}$ is the effective mass of the $\Delta$ isobar (one may put for simplicity $m^{*}_{\Delta}-m^{*}_{N}\simeq m_{\Delta}-m_{N}=293$ MeV), $\Theta(x)$ is the step function. It is taken into account that in case of the developed condensate $\mu_{n}\simeq\mu_{p}$. For more details the reader may consult Refs. Voskresensky et al. (1977); Migdal et al. (1990). The electric charge of protons is screened by the condensate of negative pions and by electrons and muons produced via the weak $\beta$ processes and from the vacuum (at rather large values of ZZVoskresensky et al. (1977). The electric potential is found from the Poisson equation for a spherical system of radius $R$ and constant baryon density $n$,
$\Delta V$ $=4\pi e^{2}(n_{e}+n_{\mu}+n_{h})\,$,
@@
Here $f^{2}_{\pi}V$ is the contribution of the developed $\pi^{-}$ condensate as it follows from the consideration within the $\sigma$ model and $n/2$ is the total baryon charge density in the presence of the developed condensate Baym et al. (1975), $n_{e}$ and $n_{\mu}$ are contributions of electrons and muons. Nullifying the r.h.s. of Eq. (29), similarly to that we did above for Eq. (19), demonstrates the charge neutrality condition in the interior of a large size system.
@@
Figure 5: Binding energies per nucleon for the PNM (solid lines), the ISM (dashed lines), and for the beta-equilibrium matter ($pn\Delta$) with the developed $\pi^{-}$ condensate (dash-dotted lines). The latter lines are calculated for several values of the Landau-Migdal parameter $g^{\prime}$ characterising the strength of the short-range nucleon-nucleon correlations; the parameters are shown by labels on the lines. On the left panel, the nuclear part of the energy is taken from the FP equation of state Friedman and Pandharipande (1981). On the right panel we use the RMF model Cubero et al. (1988). Extra explanations are given in the paper body. The choice of parameters is such that the RMF equation of state supports the maximum neutron star mass of $2.09\,M_{\odot}$ Kolomeitsev and Voskresensky (2005).

### v2

@@ III.4 Anomalous supercharged nuclei consisting dark matter
In 1981 Cahn and Glashow [48] assumed existence of superheavy charged dark matter particles, cf. also [63]. In 2005 Glashow [103] conjectured that more specifically the dark matter may consist of the U tera-quarks and O-UUU (charge 2-2) baryons, and in 2006 Khlopov [142] studied OOHe-hidden dark matter with $m_{O}\sim\!1$ TeV. Reference [95] suggested existence of stable $O$-nuclearites of arbitrary size (up to $A\sim 10^{57}$), if there exist charged dark matter stable superheavy particles. In this case the proton charge of the nucleus is compensated by the heavy $O$-particles. Due to the large mass of the $O$ particles they yield a negligible contribution to the kinetic energy. Therefore, the ordinary ISM at density $n\simeq n_{0}$ with embedded $O$-particles compensating the charge would be absolutely stable. Similarly to estimates (17), (20) performed above for supercharged nuclei, in case of the nuclear–$O$ matter one may present the energy as
$\mathcal{E}\simeq\mathcal{E}_{0}A-\int d^{3}x(n_{p}-2n_{O})V\,.$ (24)
@@
At present there exist various anomalous events, some of which may have a relation to the discussed above scalar fields, $\pi^{-}$ condensate, quark or dark matter nuggets, see [301] and references therein and a discussion in Sect. VIII below.
IV Pion mode and condensation in dense nuclear systems
@@
The phenomenon of a pion condensation was proposed in Refs. [204, 205, 246, 244] and investigated then in numerous works, see [215, 210, 284] and references therein. The pion-nucleon interaction is strongly attractive in the p wave. Therefore, starting from the first works one expected that a pion condensate is formed in a state with a finite momentum $k\neq 0$ when the nucleon density exceeds some critical density $n_{c}^{\rm(p)}$ [213, 206, 23, 208, 49, 50, 22, 62, 265]. In first works one anticipated that $n_{c}^{\rm(p)}$ might be smaller than $n_{0}$. Soon, after analyses of atomic nucleus data, it became clear that $n_{c}^{\rm(p)}>n_{0}$ because of a stronger repulsive effect of short-range nucleon-nucleon correlations, see  [215, 210] and references therein. In Ref. [214] Migdal used the name “superheavy anomalous nuclei” and then “superdense” nuclei in Ref. [215] and later in Ref. [210] simply “abnormal” pion condensate nuclei. These objects are neutron-rich to reduce the repulsive Coulomb energy. The key idea is that for $n>n_{c}^{\rm(p)}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}n_{0}$ there may appear the $\pi^{+}$, $(\pi^{+}\pi^{-})$, or $\pi^{0}$ condensates leading a decrease in the energy density by a value which can be roughly estimated as
$E_{\pi}\simeq-\frac{1}{2}\beta_{0}(n)(n-n_{c}^{\rm(p)})^{2}\theta(n-n_{c})$
with $\beta_{0}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1/m_{\pi}^{2}$, compare with Eq. (13) that was used above in the case of the Lee-Wick $\sigma$ condensate. The quantity $\beta_{0}$ depends on the value of the Landau-Migdal parameter $g^{\prime}$, which is not well determined for $n\neq n_{0}$ and for isospin-asymmetric matter. For a smaller $\beta_{0}$, the pion condensate only slightly modifies the total energy density. For a larger $\beta_{0}$ the gain in the condensate energy may overwhelm the loss in the purely baryon repulsive contribution. In the latter case there could exist metastable or even absolutely stable abnormal $\pi$ condensate nuclei and nuclei-stars.
IV.1 Pion-nucleon interaction amplitude and pion condensation
Resummation of a nucleon-nucleon interaction in the particle–hole channel in the Fermi liquid approach yields the following representation of the $NN$ interaction amplitude in the spin-isospin channel [215, 210]
[…]
Here $\tau$ and $\sigma$ are Pauli matrices acting in the isospin and spin spaces, respectively, $g^{\prime}$ is the Landau-Migdal parameter of $NN$ correlations in the spin-isospin channel, $C_{0}^{-1}$ is the normalization factor – the density of states at the Fermi surface at $n=n_{0}$. Similar equation holds for separated $nn(pp)$ and $np$ spin channels in isospin-asymmetric matter, where $g_{nn(pp)}$ and $g_{np}$ Landau-Migdal parameters enter instead of the $g^{\prime}$ parameter. Note that the $\Delta$-nucleon Landau-Migdal parameter $g^{\prime}_{\Delta N}\ll g^{\prime}$ and thereby we for simplicity disregarded this contribution in (25), whereas the $\Delta$ contribution in the second term of Eq. (25) is taken into account, see [210] for a more detail. $\Phi(\omega,k,n)$ stands for the famous Lindhard function corresponding to the particle-hole loop
[…]
For nonzero temperature $T$ it also depends on $T$. The first term in r.h.s. of Eq. (25) can be cast as the $NN$ vertex correlation factor
$\Gamma(x,\omega,k,n)=\frac{1}{1+2\,x\,\frac{p_{\rm F}(n)}{p_{\rm F}(n_{0})}\Phi(\omega,k,n)}\,$, (27)
which appears not only in the spin-isopsin channel (for $x=g^{\prime}$) but in other channels with other Landau–Migdal parameters $x=f$, $f^{\prime}$, $g$. The wavy bold line represents the full pion Green function. For $\omega\ll m_{\pi}$ and $k\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}p_{\rm F}$ the main contribution to the $NN$ amplitude is determined by the second diagram in r.h.s. of (25).
@@ IV.1 Pion-nucleon interaction amplitude and pion condensation
Main contributions to the pion polarization operator come from the nucleon–nucleon-hole and $\Delta$–nucleon-hole terms corrected by the short-range correlations, for more details see [210, 284, 288]. The contribution of the $\Delta$–nucleon-hole is as important as that of the nucleon-nucleon hole. Indeed, the mass of the $\Delta$ isobar is only $2m_{\pi}$ larger than the nucleon mass but the degeneracy factor is four times larger and the attractive $\pi N\Delta$ coupling constant is twice as large as the $\pi NN$ one.
The retarded pion propagator is
@@
$-[D_{\pi}^{R}(\mu_{\pi},\vec{k})]^{-1}$ $=\widetilde{\omega}^{2}(\mu_{\pi},k^{2})$
$=m^{2}_{\pi}+k^{2}+\Re\Pi^{R}(\mu_{\pi},k^{2})-\mu^{2}_{\pi}\,$, (29)
$\Re\Pi(\omega,k)$ is the real part of the pion polarization operator, the imaginary part $\Im\Pi(\mu_{\pi},k)=0$, $f_{\pi NN}\approx m_{\pi}^{-1}$ is the $\pi NN$ coupling constant, and $\Gamma(n)$ is the factor taking into account the $NN$ correlations, $\Gamma(n\to 0)\to 1$, $\mu_{\pi}$ is the pion chemical potential. In the case of the ISM, $\mu_{\pi^{\pm}}=0$ in accordance with reactions $p\leftrightarrow n+\pi^{+}$, $n\leftrightarrow p+\pi^{-}$; $\mu_{\pi^{0}}=0$ for the case of arbitrary isospin composition in accordance with reactions $N\leftrightarrow N+\pi^{0}$, $N=(n,p)$. The quantity
$\widetilde{\omega}^{2}_{0}=\min_{k}\{\widetilde{\omega}^{2}(\mu_{\pi},k)\}=\widetilde{\omega}^{2}(k_{0})\,$, (30)
where the minimum is realized at a density dependent momentum $k_{0}(n)$, has the meaning of the squared effective pion gap, see Ref. [210] and references therein. We notice that the in-medium $\pi^{-}$ can play a role of the light negatively charged boson considered in Sect. III with the parameter $m_{b}^{2}$ used there replaced by the squared effective pion gap $\widetilde{\omega}^{2}_{0}$.
At densities $n\sim n_{c}^{\rm(p)}$ the pion gap behaves like $\widetilde{\omega}^{2}_{0}\sim\beta(n_{c}^{\rm(p)}-n)$, with a constant $\beta(\mu_{\pi},k_{0})>0$ (at the neglect of the quantum fluctuations) and the change in the sign of the gap indicates the onset of instability with respect to the pion condensation with the momentum $k_{0}$. The critical point $n_{c}^{\rm(p)}$ is the same for $\pi^{\pm},\pi^{0}$ in ISM. In the neutron star matter, $n_{c}^{\rm(p)}=n_{c}^{\pm}$ for charged pions, and it is equal to $n_{c}^{0}$ for $\pi^{0}$. Which value $n_{c}^{\pm}$ or $n_{c}^{0}$ is larger depends on the model under consideration, see [215, 210].
In the ISM the momentum dependence of the pion gap, $\widetilde{\omega}^{2}(0,k^{2})$ develops minimum at $k=k_{0}\neq 0$ at densities $n>n_{c}^{(1)}\simeq(0.5-0.7)\,n_{0}$, see [284] and in its vicinity one may use parametriztion:
@@
where $\gamma\sim 1$. The squared effective pion gap $\widetilde{\omega}^{2}_{0}$ changes sign at $n=n_{c}^{\rm(p)}$. Considering the pion condensation in a finite system for $n>n_{c}^{\rm(p)}>n_{0}$, one may find the spatial structure of the classical pion field solving the equation
@@
which replaces Eq. (18). Nuclear systems obeying the roton-like spectra have been extensively studied, see [274, 210, 288] and references therein. From a recent time, some authors became to name such a spectrum the “moat” spectrum.
Figure 6: Typical density dependence of the squared effective pion gap for $\pi^{\pm,0}$ in ISM and for $\pi^{0}$ in neutron star matter. Solid line shows $\widetilde{\omega}^{2}(k_{0}(n))$ for $n_{c}^{(1)}<n<n_{c}^{\pi}$. For $n>n_{c}^{\pi}$ it is continued for the case when the pion condensate is not formed and the system is in a metastable state. Dotted line corresponds to the case of a saturation of the pion softenning due to artificially strong increase of the Landau–Migdal parameter $g^{\prime}$ with a density growth. Dash-dotted line characterizes the amplitude of the pion condensate appearing in the first order phase transition at $n_{c}^{\pi}$. Dashed line shows the squared effective pion gap on top of the pion condensate.
A typical density dependence of the squared effective pion gap, $\widetilde{\omega}^{2}(k_{0}(n))$, for $\pi^{\pm},\pi^{0}$ in ISM and for $\pi^{0}$ at any $N/Z$ is illustrated in Fig. 6, see also Ref. [210]. For $n>n_{c}^{(1)}$ the effective pion gap acquires a minimum at a finite momentum $k_{0}$. With a subsequent density increase, $\tilde{\omega}^{2}(k_{0}(n))$ decreases (solid line). Various calculations done within a mean-field approximation (ignoring contributions of pion fluctuations to the pion self-energy) demonstrate that at $n>n_{c}^{\pi}$ the pion condensate with a liquid-crystal-like or a solid-like structure may appear via the second-order phase transition. Actually at $n=n_{c}^{\pi}$ the classical pion field is developed, in the first-order phase transition due to the effect of fluctuations with $k\sim k_{0}\sim p_{{\rm F},N}$. Dashed-dotted line illustrates the sudden jump and the following growth of the pion condensate amplitude. The squared effective pion gap on top of the condensate is shown by the dashed line. It is growing with an increasing density. As we have noted, in the ISM the value $n_{c}^{\pi}$ is the same for $\pi^{\pm}$ and $\pi^{0}$. In the isospin-asymmetric matter the value $n_{c}^{\pi}$ depends on the sort of the pion. Actual values of the critical densities depend on poorly known density dependence of the Landau-Migdal parameters. Numerical analysis [5] performed within the “variational theory of nuclear matter”, which employs the realistic two-nucleon potential in vacuum and then introduces three-nucleon interactions, gave the value $n_{c}^{\pi}\simeq 2n_{0}$ for the critical density of the $\pi^{+},\pi^{-},\pi^{0}$ condensation in the ISM and the value $n_{c}^{\pi}\simeq 1.3n_{0}$ for the $\pi^{0}$ condensation in the purely neutron matter. Cooling of neutron stars is appropriately described within the “nuclear medium cooling scenario” which takes into account of effects of the pion softening [277, 278, 279, 210] provided $n_{c}^{\pi}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(2-2.5)n_{0}$ for $\pi^{\pm}$ and $\pi^{0}$ condensates, cf. [247, 292, 33, 110, 29, 31, 108, 109] and references therein. Continuation of the solid line for $n>n_{c}^{\pi}$ corresponds either to the possibility of a metastable phase in the system, when the ground state contains no liquid-crystal-like or solid-like pion condensates. Dotted line corresponds to the case of saturation of the pion softening because of an artificially strong increase of the Landau–Migdal parameter $g^{\prime}$ with the density.
Finally, we note that the resummation of the in-medium $NN$ interaction (25) imposes that at $n>n_{0}$ the $NN$ cross section has strong density dependence owing to the interplay of a moderately suppressing effect from the vertex correction (27) because of $NN$ correlations and a strong enhancing effect of the softening of the pion mode with the increasing density and, therefore,
$\sigma_{NN}\propto\frac{\Gamma^{4}(g^{\prime},0,k_{0},n)}{(\tilde{\omega}(k_{0}(n))/m_{\pi})^{4}}.$ (31)
IV.2 $\pi^{\pm}$ condensate nuclear systems
Spectra of pions in the ISM and purely neutron matter were demonstrated in [208]. In the latter case for $\pi^{\pm}$ there are the $\pi^{+}$ and $\pi^{-}$ branches, $\omega^{\pi^{+}}(k)$, $\omega^{\pi^{-}}(k)$ (corresponding to $\omega\to m_{\pi}$ for $n\to 0$) and the higher lying $\Delta$ branches. It proves to be that up to large densities $\omega^{\pi^{-}}(k)>\mu_{e}$ [213, 208] and the reaction $N+e\to N+\pi^{-}+\nu$ is forbidden. If pion were free, $\omega^{2}=m^{2}_{\pi}+k^{2}$, such a reaction were allowed for $\mu_{e}(n)>m_{\pi}$ resulting in efficient cooling of the neutron stars heaving such a densities in their interiors. For $n>n_{c}^{+}$ ($n_{c}^{+}<n_{0}$ according to evaluations of [208]) there appears extra branch with $\omega^{\pi^{+}}_{s}(k)<0$ corresponding to appearance of the so called $\pi^{+}_{s}$ condensate. We have mentioned this possibility in Sect. III.3. Accordingly, for $n>n_{c}^{+}$ the reaction $p\to n+\pi^{+}$ becomes possible resulting in the production of $\pi^{+}$’s occupying this branch. For $n=n_{c}^{\pm}>n_{c}^{+}$ the branches $\omega^{\pi^{+}}_{s}(k)$ and $\omega^{\pi^{-}}(k)$ coincide at $k=k_{0}\neq 0$, where $\omega^{\pi^{+}}_{s}(k_{0})+\omega^{-}_{\pi}(k_{0})=0$. Thereby for $n>n_{c}^{\pm}$ the reactions $N\leftrightarrow N+\pi^{+}+\pi^{-}$ become energetically favorable and there appears $\pi^{\pm}$ condensate. Appearance of $\pi^{\pm}$ condensation for $n>n_{c}^{\pm}$ results in the efficient neutrino cooling processes, e.g. via $N+\pi_{cond}^{-}\to N+e+\bar{\nu}$ reaction.
@@ IV.2 $\pi^{\pm}$ condensate nuclear systems
At large densities the baryon subsystem is essentially rearranged. In the approach exploiting the chiral symmetry  [49, 50, 22] it was shown that instead of two different Fermi seas for protons and neutrons, a common Fermi sea for baryon quasiparticles, being superpositions of proton neutron and $\Delta$ isobar states mixed due to interactions with the developed pion condensate, is formed. At densities much larger than $n_{c}^{\rm(p)}$ the condensate becomes developed and the squared amplitude of the pion condensate reaches its maximum value $f_{\pi}^{2}/2$ for $\pi^{\pm}$, where $f_{\pi}\approx 92$ MeV is the pion weak decay constant and the electric charge density of the baryon subsystem becomes equal to $n/2$. Taking into account these effects, and electron and muon populations, the energy density acquires the form [280]
${E}_{B+\pi+e+\mu}={E}_{B}+{E}_{\pi}+{E}_{e}+{E}_{\mu}\,$, (32)
@@
${E}_{\pi}=\frac{V^{2}f^{2}_{\pi}}{2}+\frac{1}{3}(m^{*}_{\Delta}-m^{*}_{N})\,n-\frac{81}{50}f^{2}_{\pi NN}(1-g^{\prime})n^{2}\,$,
where $E_{B}$ is the energy density of the baryonic subsystem, $g^{\prime}$ is the isospin Landau-Migdal parameter, and $m_{\Delta}^{*}$ is the effective mass of the $\Delta$ isobar (one may put for simplicity $m^{*}_{\Delta}-m^{*}_{N}\simeq m_{\Delta}-m_{N}=293$ MeV), $\Theta(x)$ is the step function. It is taken into account that in case of the developed condensate $\mu_{n}\simeq\mu_{p}$. For more details the reader may consult Refs. [280, 210]. The electric charge of protons is screened by the condensate of negative pions and by electrons and muons produced via the weak $\beta$ processes and from the vacuum (at rather large values of ZZ[280]. The electric potential is found from the Poisson equation for a spherical system of radius $R$ and constant baryon density $n$,
$\Delta V$ $=4\pi e^{2}(n_{e}+n_{\mu}+n_{h})\,$, (33)
@@
Here $f^{2}_{\pi}V$ is the contribution of the developed $\pi^{-}$ condensate as it follows from the consideration within the $\sigma$ model and $n/2$ is the total baryon charge density in the presence of the developed condensate [22], $n_{e}$ and $n_{\mu}$ are contributions of electrons and muons. Nullifying the r.h.s. of Eq. (33), similarly to that we did above for Eq. (19), demonstrates the charge neutrality condition in the interior of a large size system, for $R\ll l_{V}$ where $l_{V}$ is the Debye screening length.
@@
Figure 7: Binding energies per nucleon for the PNM (solid lines), the ISM (dashed lines), and for the beta-equilibrium matter ($pn\Delta e\mu$) with the developed $\pi^{-}$ condensate (dash-dotted lines). The latter lines are calculated for several values of the Landau-Migdal parameter $g^{\prime}$ characterising the strength of the short-range nucleon-nucleon correlations; the parameters are shown by labels on the lines. On the left panel, the nuclear part of the energy is taken from the FP equation of state [92]. On the right panel we use the RMF model [59]. Extra explanations are given in the paper body. The choice of parameters is such that the RMF equation of state supports the maximum neutron star mass of $2.09\,M_{\odot}$ [154].